Fermi National Laboratory


Accelerator Update

Friday, December 20
A safety system power supply problem caused the loss of store #2090. Safety technicians spent the rest of the day performing safety system tests. After the tests were completed, Operations established store #2091.

Saturday December 21
The Tevatron (TEV) dampers failed resulting in beam instabilities. Operations turn the dampers off for the weekend.

Sunday December 22
The Tevatron Electron Lens (TEL) failed. Experts accessed the tunnel and repaired the TEL. After resolving some Booster problems, Operations established store #2094 with an initial luminosity of 32.7E30. Recycler took the remaining antiprotons. Later, store #2094 was lost due to a kicker prefire.

Monday December 23
TeV experts conducted studies. Operations established store #2097 with an initial luminosity of 29.5E30 and Recycler took the remaining antiprotons.

Tuesday December 24
Operations terminated store #2097 that morning. After resolving some coalescing problems, Operations established store #2099 with an initial luminosity of 26.72E30.

Wednesday December 25
Operation terminated store #2099 early in the morning. Linac RF problems delayed the store #2103, which was eventually established with an initial luminosity of 32.9E30. More problems developed with Linac RF, plus an Accumulator kicker, database communications, and the Antiproton Source's (Pbar) lithium lens all contributed to poor stacking for much of the day.

Thursday December 26
Stacking resumed during the midnight shift. The Main Injector suffered two vacuum bursts, one in its RF section and one in its P1 transfer line. Operations established store #2105 with an initial luminosity of 23.8E30.

Friday December 27
Recycler experts took 30mA of antiprotons from Pbar. Unfortunately store #2105 was lost due to a kicker's power supply tripping off, which caused a quench.

Saturday December 28
While trying to remove the halo from the next store, the TeV suffered a multihouse quench due to a kicker prefire. Cryo and vacuum recovery took the rest of the day.

Sunday December 29
Cryo recovery continued. Operations established store #2113 with an initial luminosity of 31.58E30. Recycler took the remaining antiprotons.

Monday December 30
Store #2113 was intentionally terminated but the TeV quenched at sector F4 due to a problem with the antiproton orbit. Operations established store #2114.

Tuesday December 31
Operations terminated store #2114 and again the TeV quenched at F4 for the same reason as store #2113. During the next shot setup the TeV quenched again losing the store.

Wednesday January 1, 2003
Operations established store #2121 and then later established store #2123. There were problems with Main Injector RF stations (MIRF9 and MIRF1). Recycler took the remaining antiprotons.

Thursday January 2
TEL conducted end-of-store studies before Operations terminated store #2123. Technicians made repairs to TeV RF stations and Booster injection equipment.

Friday January 3
Operations established store #2134 with an initial luminosity of 28.93E30. Recycler took the remaining antiprotons. Main Injector experts conducted studies.

Saturday January 4
Operation established store #2136. Booster extraction problems delayed stacking.

Sunday January 5
Operations terminated store #2136 and established store #2138 with an initial luminosity of 29.97E30. A Linac RF station (KRF1) failed and took four hours for techs to complete its repair.

Monday January 6
The midnight shift began with experts conducting end-of-store studies. Another Linac RF station (LRF3) failed. Since the Monday day shift had been planned as a maintenance day, the station's repair was put off until the day shift. Operations terminated store #2138 at 6:49 AM. Operators turned off all the accelerators for the shutdown.

Today's Plan
The accelerators will be open for maintenance until 4 PM today. Experts will then conduct five shifts of accelerator studies.

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/6/2003   email Fermilab